Problem 14.28 (additional/static) Keebock, a maker of outstanding running shoes, keeps the soles of its size-13 running shoes in inventory for one period at a cost of $0.25 per unit. The setup costs are $50. Beginning inventory is zero and lead time is 1 week. Shown below are the net requirements per period. Period 1 2 4 5 910 Net requirements 35 30 45 010 40 30 030 55 a) Keebock's cost based on EO is (round your response to two decimal places) (Note: For the EOQ computation, the reorder point is equal to the demand during lead time, i.e., Units demanded Number of periods where units demanded is the sum of the net requirements and the number of periods is 10.)
where units demanded is the sum of the net requirements and the number of periods is 10) b) Keebock's cost based on Lot-for-lot is s(round your response to two decimal places).
Answer #1

Total demand over 10 periods = 35+30+45+0+10+40+30+0+30+55 = 275

Average demand per period = 275/10 = 27.5

EOQ = √(2*50*27.5/0.25) = 105 (rounded off)

Reorder point = Units demanded / number of periods = 275/10 = 27.5

MRP based on EOQ is following

X 0 1 2 3 4 5 6 7 8 9 10
Gross Requirements 0 35 30 45 0 10 40 30 0 30 55
Scheduled Receipts 0 0 0 0 0 0 0 0 0 0 0
Projected on Hand 0